A stockbroker trades stock for 100 days. On day one he loses 100$.Everyday after that he loses 100$ more than on the previous day. What is his average daily loss during his 100 day period? I would like to just have the formula so that I can solve it. I'm just starting alg. again in school after 27 yrs. Just need a few hints. Thanks!!!
Found 2 solutions by scott8148, edjones:
Answer by scott8148(5891)

not much of a broker

the loss for any day d is d(100), so on the hundredth day he looses 100(100) or 10000

since the loss is at a uniform rate (day-to-day change is constant); the average is (first day+last day)/2

Answer by edjones(7311)

Sum of an arithemetic sequence is: n/2(a[1]+a[n]) n=number of terms
100/2(100+10000)=50*10100=505000
Now we divide that number by n to get the average=$5,050
Ed